New vidyo room:

News:

Flavio asked for a volunteer for the VQR commission.

Framework & production

No news.

Analysis

B physics

Enrico: working on the correlation and will rerun the correlation. ETA: ~ 1 week. Martino: spoke with conveners of HFLAV. There are troubles with combination Alberto/Enrico are working on w/ Atlas colleagues, the combination is handled by the authors themselves, usually that was handled by the HFLAV. Martino: A good treatment of the correlations must be done by the authors.

Heavy composite neutrinos

MAtteo: finalizing the thesis and the answers for the preapproval. Need to add weights for the pt of the DY in the MC. Some reprocessing needed and done. Will unblind up to 1 TeV and check the weights. By next week have the note updated and get preapproved.

VBS

Matteo finished the plots needed for the thesis.

WH_SS

Activities: leptonic VH channel, providing checks on the note. Ele SF measurement: a conveners was interested in using the code for BDT application. Thesis: writing. Consolidating the unblinding results, Some plots are missing, once ready will discuss with Ugo/Alberto. First 2 chapter written. Alberto: rocco ardino joined the group and started a new analysis. master thesis (Rocco, Alberto, Jacopo, Gaia). Z/H->Jpsi gamma

Same sign dileptons

Ugo: there is some tension wrt the SM processes. Tom: atlas published a measurement of ttH where some SSdilepton was increased by 60% wrt the vanilla prediction. Some theorists suggest several models.

Htomumu/Htocharm

next week.

EPR

Detector

Trigger

DT meeting presentation on the OBDT work. Reported the observation of duplicated hits. signals coming from the firmware. Cristina proposed to have a meeting with Alvaro.

DT

most of the meeting was devoted to the Legnaro activities. Starting now the plan for the new run at the gif. Ignacio would prefer a different approach. Next week Fabrizio will give a report.

Timing Layer

This Friday there is a workshop on the peltier. Now everyone seems to be convinced that we need peltier to cool below the CO2 (which is -30C), this is to limit the dark current in the sipm. Drawback: loose 2-3% of acceptance because the crystals have to be shorter. They cost. There is a completely open area: how to power them. Make all the tests. Sandro: there is an issue. Introducing the peltier must be running all the time because you reduce the cooling power of the plate. If they are off they act as insulators. Mech eng went through the interview. HE was the only candidate.

Tracker/tracking

Study based on ML approach for cleaning doublets to decrease the seeds from pixel hits. Using the standard CPU is > 1s/ev with a max of 6000 inferences per event due to memory. Impossible. Trying to use the approach for the triplets to make it possible.

A.O.B.

Notte della ricerca: virtual visit 15:30-16:30 with many sites. the guides will be Lucia and Sonia. RobertoC will join the discussion as well. Send the info to Alberto for the website. Need a volunteer for the masterclass. -- RobertoRossin - 2020-11-17
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2020-11-17 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback